# Differentiable Fuzzy Cosmic-Web for Field Level Inference

### v1

@@ 3.2 Numerical Results of TEST1
The first numerical test is performed on a grid with $128^{3}$ voxels within a comoving volume of 1280 $h^{-1}\,\mathrm{Mpc}$ per side, corresponding to a spatial resolution of 10 $h^{-1}\,\mathrm{Mpc}$. We use a $\Phi$-web-dependent bias modeled with a power-law and negative binomial likelihood for each region, for a total of 8 bias parameters. Using this model and a fixed random seed, we construct mock catalog of object number counts, $\boldsymbol{n}^{*}$, in redshift space, which we define as the ground-truth. The priors of the power-law ($\alpha$) and negative-binomial ($\beta$) bias parameters are set to be log-normal distributed (see Eq 4) with fixed hyperparameters: $\mu_{\alpha}=0.27,\sigma_{\alpha}=0.30,\mu_{\beta}=2.12,\sigma_{\beta}=0.20$.
@@
The HMC chain is initialized with an state $(\boldsymbol{\nu}^{(0)},\boldsymbol{b}^{(0)})$, with $\boldsymbol{\nu}^{(0)}$ being the white-noise representation of the density contrast coming from Wiener-filtering $\boldsymbol{n}^{*}$, and $\boldsymbol{b}^{(0)}$ being bias parameters randomly sampled from their priors. During burn-in, chain convergence was achieved after approximately 250 samples. The adapted step-size was $\sim\!10^{-3}$ and the NUTS trajectory length adaptation saturated at 1024 leapfrog steps (corresponding to a maximum three depth of 10).
@@
After convergence, we assess the correlation length in the white-noise samples, based on their autocorrelation at lag $m$:
$\xi_{i}(m)=\frac{1}{\sigma^{2}_{i}\left(M-n\right)}\sum_{j=0}^{M-m}\left(\nu_{i}^{(j)}-\bar{\nu}_{i}\right)\left(\nu_{i}^{(j+m)}-\bar{\nu}_{i}\right)\ ,$ (18)
@@
where $\nu_{i}^{(j)}$ is the $i$th voxel of the $j$th sample, $M$ is the total number of samples of a given chain run, $\bar{\nu}_{i}$ is the within-chain mean and $\sigma_{i}^{2}$ the within-chain variance. For a chain of $M=1000$ samples, we computed $\xi_{i}(n)$ for lags $n=0$ to $100$ across $10^{4}$ randomly selected voxels. Fig. 3 shows the mean autocorrelation over this subset.
@@
Figure 3: TEST1: Mean autocorrelation $\xi(n)$ as a function of lag $n$ for a chain of length $M\approx 1000$, evaluated over $10^{4}$ randomly selected voxels. The red curve shows the ensemble average $\langle\xi(n)\rangle$ across parameters, with the shaded band indicating the ±$\pm 1\sigma$ dispersion. The horizontal dashed line marks the threshold $\xi=0.1$ used to assess mixing efficiency.
@@
We define the correlation length as the smallest lag $n$ such that $\langle\xi(n)\rangle\leq 0.1$, yielding an estimate of $\sim\!40$ samples. Given this autocorrelation we drew 15000 post-burn-in states, providing about 500 effectively independent posterior samples for the subsequent statistical analysis.
@@
We assess the computing times of single gradient calculations across different mesh sizes and bias configurations finding that they are below one second (see Fig. 4).
@@
Figure 4: Computing times for single gradient evaluations of the forward model employing ALPT evolution and the HICOBIAN bias model for different mesh sizes, run on a single NVIDIA A100-SXM4 GPU equipped with 40 GB of on-board HBM2 memory.
@@
In addition, we perform the Gelman–Rubin convergence diagnostic, which confirms that the drawn samples are consistent across chains and effectively independent (see Fig. 5).
@@
Figure 5: Gelman-Rubin convergence diagnostic computed for three independently initialized chains of 250 samples after convergence. This statistic is calculated for all cells in Fourier space, and we show their distribution with scale $k$. Values close to one indicate that the Markov chains have converged and that the samples are effectively independent.
@@
As a first step, we conduct a visual assessment of the true and reconstructed maps of the initial and final density fields (see Fig. 6). The reconstructed samples exhibit a high degree of visual similarity to the ground-truth for both the initial and final conditions. On a more quantitative level, we find that the standard deviations of the reconstructed fields display structured patterns at amplitudes approximately an order of magnitude lower than the true density field at the initial conditions. In contrast, the final conditions exhibit higher variance, reflecting the increased nonlinearity and complexity of the evolved structures.
@@
To quantitatively assess the quality of the reconstruction, we compute both two- and three-point statistics in Fourier space for the initial (Fig. 7) and final conditions (Fig. 8). Specifically, we evaluate the monopole and quadrupole moments of the power spectrum, along with the reduced bispectrum for a configuration particularly sensitive to nonlinear effects ($k_{1}=0.1$ and $k_{2}=0.2$ $h\,\mathrm{Mpc}^{-1}$). In addition, we compute the propagator, defined as the normalized cross-correlation in Fourier-space between the ground-truth and the reconstructed initial density field. The motivation for investigating the quadrupole and bispectrum of a the reconstructed Gaussian field is to assess how well the samples recover deviations introduced by cosmic variance in these statistics.
@@
It is remarkable to observe in Figs. 7 and 8 that the monopoles, quadrupoles and bispectra are accurately recovered not only in the final conditions but also in the initial conditions, even reproducing the specific features inherent to the particular cosmic variance realization. Although local bias models fail to reproduce three-point statistics without short- and long-range nonlocal terms (Coloma-Nadal et al., 2024), we show that neglecting long-range nonlocal bias only also leads to significant differences in the bispectrum (see third panel in Fig. 8).
@@
The propagator shows a significant gain in information, demonstrating the effectiveness of the reconstruction. To evaluate how close the result is to an ideal scenario, we compute the optimal cross at final conditions by generating a synthetic tracer number count field from the evolved ground-truth dark matter field, using the exact same bias parameters and stochastic prescription as in the model, but a different seed. The cross-correlation between this ideal tracer field and the ground-truth serves as a benchmark, represented by the dashed-dotted red line in the figure. This allows us to assess whether the reconstruction approaches the theoretical limit given the assumed bias model. As we find in Fig. 8, the red and blue lines of the right panel are on top of each other, indicating that the optimal cross correlation limited by shot noise has been reached.
@@
In Fig. 9 we show a corner plot of the posterior distributions of the eight parameters of the bias model compared to their values in the reference catalog. The power-law parameters $\alpha$ are recovered with sub-percent accuracy, while the negative-binomial $\beta$-parameter is recovered to few percent precision. The absence of significant correlations between $\alpha$ bias parameters in distinct cosmic-web regions supports the validity and stability of our inference framework.
@@
Overall, the consistency of all diagnostics—maps, summary statistics, and bias-parameter posteriors—demonstrates that the complete field-level inference pipeline reliably recovers both the underlying density field and the cosmic-web-dependent bias parameters at the targeted resolution. The numerical studies TEST2 and TEST3 further confirm the validation of the BRIDGE code (details can be found in Appendix C).
@@
Figure 6: TEST1: Spatial summary of 500 independent reconstructions ($N=128$). Slices of $\Delta L=10\,h^{-1}\mathrm{Mpc}$ (one voxel width). Top row: reconstructed initial density contrast. Bottom row: reconstructed tracer field. Columns from left to right: (1) the pixel-wise standard deviation across the 500 independent reconstructions, (2) the mean reconstructed field, (3) one representative reconstruction sample, and (4) the true reference field.
@@
Figure 7: TEST1: Comparison of the summary statistics of the initial density contrast field for the mock reference (red dashed) and 500 independent reconstructions (blue solid). Columns from left to right: (1) the monopole power spectrum, (2) the quadrupole, (3) the reduced bispectrum as a function of triangle opening angle $\theta$, and (4) the cross-correlation coefficient $C(k)$ between the reference and reconstruction (blue solid) and the propagator (dashed red), defined as the cross-correlation between the mock tracer density contrast and the initial density reference. In columns (1)–(3), the lower sub-panels display the ratios with respect to the reference. Shaded bands denote the $1\sigma$ intervals from the posterior samples. Vertical dashed lines indicate the isotropic Nyquist frequency, $k_{\mathrm{nyq}}/\sqrt{3}$.
@@
Figure 8: TEST1: Comparison of the summary statistics of the mock reference, and 500 independent reconstructions evolved through the forward model. Panel definitions are identical to those in Figure 7, except that in the rightmost panel the propagator is replaced by the maximal tracer field cross-correlation and that we show the bispectrum of the dark-matter field and a tracer field using a local bias (without cosmic web) with bias parameters set to those of knots.
@@
Figure 9: TEST1: Posterior distributions for the 500 independent samples of the $\alpha$ and $\beta$ bias parameters in each cosmic-web region of the $\Phi$-web. Red dots and lines show the values of the parameters for the ground-truth catalogue. The posteriors are smoothed with a 2D Gaussian filter with standard deviation equal to 0.75 bin widths. TEST2 yields equivalent results.
@@ 4 Conclusions
In this work, we have developed and validated a novel Bayesian field-level inference framework that incorporates a physically motivated, differentiable, and nonlocal bias model–HICOBIAN–based on the hierarchical cosmic-web. This model allows for smooth transitions between different cosmic environments via a fuzzy classification approach, making it well-suited for machine learning integration and GPU acceleration.
We demonstrated that our approach accurately reconstructs the initial density field, achieving high fidelity even in the presence of redshift-space distortions and complex, nonlinear, nonlocal and stochastic bias. The reconstruction quality was evaluated using two- and three-point statistics, including the power spectrum monopole and quadrupole, the reduced bispectrum, and the propagator, showing excellent agreement with the ground-truth and reaching the optimal limit set by shot noise.
We further tested the robustness of the model under various conditions, including varying spatial resolutions and complex bias scenarios involving 4 to 16 cosmic-web regions, each characterized by distinct bias parameters. These included power-law bias, exponential threshold bias, and deviations from Poisson statistics, totaling 8 to 64 bias parameters. Across all settings, the method maintained its high accuracy, confirming the generality and adaptability of our approach.
In future work, we plan to investigate the performance of the BRIDGE code with high-fidelity synthetic reference catalogues and directly apply it to redshift survey data across various tracers, including Bright Galaxies (BGs), Luminous Red Galaxies (LRGs), Emission Line Galaxies (ELGs), Quasars (QSOs), and Lyman-$\alpha$ forests. This will require incorporating light-cone evolution, selection effects and conducting a more detailed analysis of redshift-space distortions in the highly nonlinear regime.
By implementing this framework in the GPU-accelerated BRIDGE code using JAX, we enable efficient and scalable inference on large datasets. This work thus marks a significant step toward a unified, interpretable, and fully differentiable framework for cosmological data analysis, with direct applicability to tracers of the large-scale structure.
@@ Appendix A White-Noise Parametrization
$\mathsf{F}^{-1}_{nm}$ $=L^{-3}\exp\left(i\boldsymbol{k}_{m}\cdot\boldsymbol{x}_{n}\right)\,,$ (21)
Here $\boldsymbol{x}_{n}$ represents the cartesian direction vector in commoving space at a voxel indexed by $n$. The same holds for the wave-vector $\boldsymbol{k}_{n}$ in the Fourier-transformed space. We have defined the diagonal linear matter power spectrum matrix $\mathsf{P}_{nm}\equiv\delta^{\mathrm{K}}_{nm}\,P_{\mathrm{lin}}(|\boldsymbol{k}_{n}|)$. The density contrast, $\boldsymbol{\delta}_{0}$, is then Gaussian-distributed with zero mean and covariance
@@
In principle, one can choose to parametrize the field either in terms of $\boldsymbol{\delta}_{0}$ directly or through the underlying white noise $\boldsymbol{\nu}$. Sampling $\boldsymbol{\delta}_{0}$ directly removes the linear transformation from the forward model but the FFTs must still be performed at every MCMC step to evaluate the negative log prior and to draw new field realisations. We have tested both parametrizations and did not observe significant differences in computational cost or sampling behavior. Nonetheless, we favor the white-noise parametrization $\boldsymbol{\nu}$ for its simpler implementation.
@@
$\boldsymbol{\delta}_{0}=\sqrt{\frac{L^{3}}{2}}\,\mathsf{F}^{-1}\sqrt{\mathsf{P}}\,\hat{\boldsymbol{u}}\ .$ (24)
With this parametrization every parameter correponds to a definite scale/wavenumber. We expected different chain sampling behavior especially when using diagonal mass matrix adaptation during burn-in, but we saw no significant differences with respect to the white noise parametrization. We note that this parametrization requires careful construction of an Hermitian-packed Fourier ($N,N,N/2+1$)-shaped array produced by a 3-D real-to-complex FFT out of $N^{3}$ random real numbers, with explicit handling of the array structure and Nyquist planes.
@@
  • $N^{3}/2-4$ phases uniformly distributed in $[0,2\pi)$
@@
  • $8$ Nyquist modes following a Gaussian distribution.
. This parametrization may be advantageous if we want to fix the amplitudes to the theoretical linear matter power spectrum, effectively reducing the parameter space in half. But again, for our work, we did not come across any obvious advantages of using this approach over the more straightforward white-noise parametrization.

### v3

@@ 3.2 Numerical results of TEST1
The first numerical test was performed on a grid with $128^{3}$ voxels within a comoving volume of 1280 $h^{-1}\,\mathrm{Mpc}$ per side, corresponding to a spatial resolution of 10 $h^{-1}\,\mathrm{Mpc}$. We used a $\Phi$-web-dependent bias modelled with a power-law and negative-binomial likelihood for each region, for a total of eight bias parameters. Using this model and a fixed random seed, we constructed a mock catalogue of object number counts, $\boldsymbol{n}^{*}$, in redshift space, which we define as the ground truth. The priors of the power-law ($\alpha$) and negative-binomial ($\beta$) bias parameters were set to be log-normally distributed (see Eq 4) with fixed hyperparameters: $\mu_{\alpha}=0.27,\sigma_{\alpha}=0.30,\mu_{\beta}=2.12,\sigma_{\beta}=0.20$.
@@
The HMC chain was initialised with an state $(\boldsymbol{\nu}^{(0)},\boldsymbol{b}^{(0)})$, with $\boldsymbol{\nu}^{(0)}$ being the white-noise representation of the density contrast coming from Wiener-filtering, $\boldsymbol{n}^{*}$, and $\boldsymbol{b}^{(0)}$ being bias parameters randomly sampled from their priors. During burn-in, chain convergence was achieved after approximately 250 samples. The adapted step size was $\sim\!10^{-3}$, and the NUTS trajectory length adaptation saturated at 1024 leapfrog steps (corresponding to a maximum tree depth of 10).
@@
After convergence, we assessed the correlation length in the white-noise samples, based on their autocorrelation at lag $m$:
$\xi_{i}(m)=\frac{1}{\sigma^{2}_{i}{\left(M-m\right)}}\sum_{j=0}^{M-m}\left(\nu_{i}^{(j)}-\bar{\nu}_{i}\right)\left(\nu_{i}^{(j+m)}-\bar{\nu}_{i}\right)\ ,$ (20)
@@
where $\nu_{i}^{(j)}$ is the $i$th voxel of the $j$th sample, $M$ the total number of samples of a given chain run, $\bar{\nu}_{i}$ the within-chain mean, and $\sigma_{i}^{2}$ the within-chain variance. For a chain of $M=1000$ samples, we computed $\xi_{i}(m)$ for lags $m=0$ to $100$ across $10^{4}$ randomly selected voxels. Figure 3 shows the mean auto-correlation over this subset.
@@
Figure 3: TEST1: Mean auto-correlation $\xi(n)$ as a function of lag, $n,$ for a chain of length of $M\approx 1000$, evaluated over $10^{4}$ randomly selected voxels. The red curve shows the ensemble average, $\langle\xi(n)\rangle,$ across parameters, with the shaded band indicating the $\pm 1\sigma$ dispersion. The horizontal dashed line marks the threshold, $\xi=0.1,$ used to assess mixing efficiency.
@@
We defined the correlation length as the smallest lag, $n,$ such that $\langle\xi(n)\rangle\leq 0.1$, yielding an estimate of $\sim\!40$ samples. Given this auto-correlation, we drew 15000 post-burn-in states, providing about 500 effectively independent posterior samples for the subsequent statistical analysis.
@@
We assessed the computing times of single-gradient calculations across different mesh sizes and bias configurations finding that they are below one second (see Fig. 4).
@@
Figure 4: Computing times for single-gradient evaluations of the forward model employing ALPT evolution and the HICOBIAN bias model for different mesh sizes, run on a single NVIDIA A100-SXM4 GPU equipped with 40 GB of on-board HBM2 memory. For reference, the full burn-in and sampling procedure for the $128^{3}$ case in this study required approximately 100 GPU hours.
@@
In addition, we performed the Gelman–Rubin convergence diagnostic, which confirms that the drawn samples have converged to the same target distribution (see Fig. 5).
@@
Figure 5: Gelman–Rubin convergence diagnostic computed for three independently initialised chains of 250 samples after convergence. This statistic is calculated for all cells in Fourier space, and we show their distribution with scale, $k$. Values close to one indicate that the drawn samples have converged to the same distribution.
@@
As a first step, we conducted a visual assessment of the true and reconstructed maps of the initial and final density fields (see Fig. 6). The reconstructed samples exhibit a high degree of visual similarity to the ground truth for both the initial and final conditions. On a more quantitative level, we find that the standard deviations of the reconstructed fields display structured patterns at amplitudes approximately an order of magnitude lower than the true density field at the initial conditions. In contrast, the final conditions exhibit higher variance, reflecting the increased non-linearity and complexity of the evolved structures.
@@
To quantitatively assess the quality of the reconstruction, we computed both two- and three-point statistics in Fourier space for the initial (Fig. 7) and final conditions (Fig. 8). Specifically, we evaluated the monopole and quadrupole moments of the power spectrum, along with the reduced bispectrum for a configuration particularly sensitive to non-linear effects ($k_{1}=0.1$ and $k_{2}=0.2$ $h\,\mathrm{Mpc}^{-1}$). In addition, we computed the propagator, defined as the cross-correlation between the mock tracer density contrast and the initial density reference. The motivation for investigating the quadrupole and bispectrum of a the reconstructed Gaussian field was to assess how well the samples recover deviations introduced by cosmic variance in these statistics.
@@
In Figs. 7 and 8, it is remarkable to observe that the monopoles, quadrupoles, and bispectra are accurately recovered not only in the final conditions, but also in the initial conditions, even reproducing the specific features inherent to the particular cosmic variance realisation. Although local-bias models fail to reproduce three-point statistics without short- and long-range non-local terms (Coloma-Nadal et al., 2024), we show that only neglecting long-range non-local bias also leads to significant differences in the bispectrum (see the third panel in Fig. 8).
@@
The propagator ensures a significant gain in information, demonstrating the effectiveness of the reconstruction. To evaluate how close the result is to an ideal scenario, we computed the optimal cross at final conditions by generating a synthetic-tracer number-count field from the evolved ground-truth dark-matter field using the exact same bias parameters and stochastic prescription as in the model, but a different seed. The cross-correlation between this ideal tracer field and the ground truth serves as a benchmark, represented by the dashed-dotted red line in the figure. This allowed us to assess whether the reconstruction approaches the theoretical limit given the assumed bias model. As we see in Fig. 8, the red and blue lines of the right panel are on top of each other, indicating that the optimal cross-correlation limited by shot noise has been reached.
@@
In Fig. 9, we show a corner plot of the posterior distributions of the eight parameters of the bias model compared to their values in the reference catalogue. The power-law parameters, $\alpha,$ were recovered with sub-percent accuracy, while the negative binomial $\beta$-parameter was recovered with precision of a few percent. The absence of significant correlations between $\alpha$ bias parameters in distinct cosmic-web regions supports the validity and stability of our inference framework.
@@
Overall, the consistency of all diagnostics – maps, summary statistics, and bias-parameter posteriors – demonstrates that the complete field-level inference pipeline reliably recovers both the underlying density field and the cosmic-web-dependent bias parameters at the targeted resolution. The numerical studies TEST2 and TEST3 further confirm the validation of the BRIDGE code (details can be found in Appendix C). However, it is remarkable that we did not recover the original bias parameters when we allowed them to vary freely in case TEST3.
@@
Figure 6: TEST1: Spatial summary of 500 independent reconstructions ($N=128$). Slices of $\Delta L=10\,h^{-1}\mathrm{Mpc}$ (one voxel width). Top row: Reconstructed initial density contrast. Bottom row: Reconstructed tracer field. Columns from left to right: (1) Pixel-wise standard deviation across the 500 independent reconstructions; (2) mean reconstructed field; (3) one representative reconstruction sample; and (4) true reference field.
@@
Figure 7: TEST1: Comparison of summary statistics of the initial density contrast field for the mock reference (dashed red line) and 500 independent reconstructions (solid blue line). Columns from left to right: (1) Monopole power spectrum; (2) quadrupole; (3) reduced bispectrum as a function of triangle opening angle, $\theta$; and (4) cross-correlation coefficient $C(k)$ between the reference and reconstruction (solid blue line) and the propagator (dashed red line), defined as the cross-correlation between the mock tracer density contrast and the initial density reference. In columns (1)–(3), the lower sub-panels display the ratios with respect to the reference. Shaded bands denote the $1\sigma$ intervals from the posterior samples. Vertical dashed lines indicate the isotropic Nyquist frequency, $k_{\mathrm{nyq}}/\sqrt{3}$.
@@
Figure 8: TEST1: Comparison of summary statistics of the mock reference and 500 independent reconstructions evolved through the forward model. Panel definitions are identical to those in Fig. 7, except that in the rightmost panel the propagator is replaced by the maximal tracer field cross-correlation and that we show the bispectrum of the dark-matter field and a tracer field using a local bias (without cosmic web) with bias parameters set to those of knots.
@@
Figure 9: TEST1: Posterior distributions for the 500 independent samples of the $\alpha$ and $\beta$ bias parameters in each cosmic-web region of the $\Phi$ web. Red dots and lines show the values of the parameters for the ground-truth catalogue. The posteriors are smoothed with a 2D Gaussian filter with standard deviation equal to 0.75 bin widths. TEST2 yields equivalent results.
@@ 4 Conclusions
In this work we developed and validated a novel Bayesian field-level inference framework that incorporates a physically motivated, differentiable, and non-local bias model – HICOBIAN – based on the hierarchical cosmic-web. This model allows for smooth transitions between different cosmic environments via a fuzzy classification approach, making it well suited for machine-learning integration and GPU acceleration.
We demonstrated that, within a self-specified setting where inference is performed on data generated by the exact same forward model, our approach accurately reconstructs the initial density field, achieving high fidelity even in the presence of redshift-space distortions and complex, non-linear, non-local, and stochastic bias. The reconstruction quality was evaluated using two- and three-point statistics, including the power-spectrum monopole and quadrupole, the reduced bispectrum, and the propagator, showing excellent agreement with the ground truth and reaching the optimal limit set by shot noise.
We further tested the robustness of the model under various conditions, including varying spatial resolutions and complex bias scenarios involving four to 16 cosmic-web regions, each characterised by distinct bias parameters. These included power-law bias, exponential threshold bias, and deviations from Poisson statistics, totalling eight to 64 bias parameters. Across all settings, the method maintained its high accuracy, confirming the generality and adaptability of our approach. However, for the complex bias models, the original bias parameters were not recovered when left to vary freely. This hints at the complexity of bias modelling and the potential limitations of field-level inference approaches that ignore intra-halo dynamics. As demonstrated in recent work (e.g. Favole et al., 2025), a realistic halo-to-galaxy connection based on high-resolution $N$-body simulations requires a large number of parameters to capture processes such as halo exclusion, satellite dynamics, quenching, and intra-halo flows.
In future work, we plan to investigate the performance of the BRIDGE code with high-fidelity synthetic reference catalogues and directly apply it to redshift survey data across various tracers, including DESI-tracers, such as bright galaxies, luminous red galaxies, emission-line galaxies, quasars, and Lyman-$\alpha$ forests. This will require the incorporation of light-cone evolution and selection effects and a more detailed analysis of redshift-space distortions in the highly non-linear regime. Moreover, our results suggest that achieving full consistency may require a joint Bayesian inference framework that incorporates intra-halo dynamics informed by detailed $N$-body simulations to break the degeneracy in the bias parameters.
By implementing this framework in the GPU-accelerated BRIDGE code using JAX, we enable efficient and scalable inference on large datasets. This work thus marks a significant step towards a unified, interpretable, and fully differentiable framework for cosmological data analysis, with direct applicability to tracers of the large-scale structure.
@@ Appendix A White-noise parametrization
$\mathsf{F}^{-1}_{nm}$ $=L^{-3}\exp\left(i\boldsymbol{k}_{m}\cdot\boldsymbol{x}_{n}\right)\,.$ (23)
Here $\boldsymbol{x}_{n}$ represents the Cartesian direction vector in commoving space at a voxel indexed by $n$. The same holds for the wave-vector $\boldsymbol{k}_{n}$ in the Fourier-transformed space. We have defined the diagonal linear matter power spectrum matrix $\mathsf{P}_{nm}\equiv\delta^{\mathrm{K}}_{nm}\,P_{\mathrm{lin}}(|\boldsymbol{k}_{n}|)$. The density contrast, $\boldsymbol{\delta}_{0}$, is then Gaussian-distributed with zero mean and covariance:
@@
In principle, one can choose to parametrize the field either in terms of $\boldsymbol{\delta}_{0}$ directly or through the underlying white noise $\boldsymbol{\nu}$. Sampling $\boldsymbol{\delta}_{0}$ directly removes the linear transformation from the forward model but the fast Fourier transforms must still be performed at every Markov chain Monte Carlo step to evaluate the negative log prior and to draw new field realisations. We have tested both parametrizations and did not observe significant differences in computational cost or sampling behaviour. Nonetheless, we favour the white-noise parametrization $\boldsymbol{\nu}$ for its simpler implementation.
@@
$\boldsymbol{\delta}_{0}=\sqrt{\frac{L^{3}}{2}}\,\mathsf{F}^{-1}\sqrt{\mathsf{P}}\,\hat{\boldsymbol{u}}\$. (26)
With this parametrization every parameter corresponds to a definite scale/wavenumber. We expected a different chain sampling behaviour, especially when using diagonal mass matrix adaptation during burn-in, but we saw no significant differences with respect to the white noise parametrization. We note that this parametrization requires the careful construction of a Hermitian-packed Fourier ($N,N,N/2+1$)-shaped array produced by a 3D real-to-complex fast Fourier transform out of $N^{3}$ random real numbers, with explicit handling of the array structure and Nyquist planes.
@@
  • $N^{3}/2-4$ phases uniformly distributed in $0,2\pi)$
@@
  • 88 Nyquist modes following a Gaussian distribution.
This parametrization may be advantageous if we want to fix the amplitudes to the theoretical linear matter power spectrum, effectively reducing the parameter space in half. But again, for our work, we did not come across any obvious advantages of using this approach over the more straightforward white-noise parametrization.